Mickley Corporation produces two products, Alpha6s and Zeta7s, which pass through two operations, Sintering and Finishing Each of the products uses two raw materials-X442 and Y661. The company uses a standard cost system, with the following standards for each product (on a per unit basis) Raw Material Standard Labor Time Y661 1.5 liters 2.5 liters Product Alpha6 Zeta7 X442 2.0 kilos 3.0 kilos Sintering 0.10 hours 0.45 hours Finishing 0.80 hours 0.90 hours Information relating to materials purchased and materials used in production during May follows Used in Production 9,300 kilos 13,800 liters Purchase Cost Standard Price Material X442 Y661 Purchases 14,800 kilos$66,600 $4.30 per kilo 15,800 liters $23,700 $1.60 per liter The following additional information is available a. The company recognizes price variances when materials are purchased b. The standard labor rate is $22.00 per hour in Sintering and $21.00 per hour in Finishing c. During May, 1,220 direct labor-hours were worked in Sintering at a total labor cost of $28,060, and 2,930 direct labor-hours were worked in Finishing at a total labor cost of $67,390 d. Production during May was 2,300 Alpha6s and 1,600 Zeta7s
Req 1 Req 2 Req 3 Complete the standard cost card for each product, showing the standard cost of direct materials and direct labor. (Round your answers to 2 decimal places.) Standard Quantity or Hours Standard Price or Rate Standaro Cost Alpha6 Direct materials-X442 Direct materials-Y661 Direct labor-Sintering Direct labor-Finishing Total kilos iters hours hours per kilo per liter per hour per hour Zeta7 Direct materials-X442 Direct materials-Y661 Direct labor-Sintering Direct labor-Finishing Total kilos liters hours hours per kilo per liter per hour per hour
Req 1 Req 2 Req 3 Compute the materials price and quantity variances for each material. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.) Direct Materials Variances-Material X442: Materials price variance Materials quantity variance Direct Materials Variances-Material Y661 Materials price variance Materials quantity variance 〈 Req 1 Req 3>
Answer #1
1 Standard cost card:
Standard quantity
or hours
Standard price
or rate
Standard
cost
Alpha6:
Direct materials-X442 2 Kilos 4.3 Per kilo 8.6
Direct materials-Y661 1.5 Liters 1.6 Per liter 2.4
Direct labor-Sintering 0.1 Hours 22 Per hour 2.2
Direct labor-Finishing 0.8 Hours 21 Per hour 16.8
Total 30
Zeta7:
Direct materials-X442 3 Kilos 4.3 Per kilo 12.9
Direct materials-Y661 2.5 Liters 1.6 Per liter 4
Direct labor-Sintering 0.45 Hours 22 Per hour 9.9
Direct labor-Finishing 0.9 Hours 21 Per hour 18.9
Total 45.7
2 Material X442:
Material price variance=Actual material purchased*(Standard rate for material-Actual rate for material)
Actual material purchased=14800 Kilos
Standard rate=$ 4.3 per kilo
Actual rate=66600/14800=$4.5 per kilo
Material price variance=14800*(4.3-4.5)=2960 U
(Actual rate is more than the standard rate.Hence, variance is unfavorable)
Material quantity variance=Standard rate for materials*(Standard materials required-Actual material used)
Standard rate=$ 4.3 per kilo
Standard materials required=Actual units produced*Material required per unit=(2300*2)+(1600*3)=9400 kilos
Actual material used=9300 kilos
Material quantity variance=4.3*(9400-9300)=430 F
(Actual material used is less than the standard material required.Hence, variance is favorable)
Material Y661:
Material price variance=Actual material purchased*(Standard rate for material-Actual rate for material)
Actual material purchased=15800 liters
Standard rate=$ 1.6 per liter
Actual rate=23700/15800=$1.5 per liter
Material price variance=15800*(1.6-1.5)=1580 F
(Actual rate is less than the standard rate.Hence, variance is favorable)
Material quantity variance=Standard rate for materials*(Standard materials required-Actual material used)
Standard rate=$ 1.6 per liter
Standard materials required=Actual units produced*Material required per unit=(2300*1.5)+(1600*2.5)=7450 liters
Actual material used=13800 liters
Material quantity variance=1.6*(7450-13800)=10160 U
(Actual material used is more than the standard material required.Hence, variance is unfavorable)
3 Sintering:
Labor rate variance=Actual hours worked*(Statndard labor rate-Actual labor rate)
Actual hours worked=1220 hours
Standard labor rate=$22 per hour
Actual labor rate=Actual direct labor cost/Actual hours=28060/1220=$23 per hour
Labor rate variance=1220*(22-23)=1220 U
(Actual labor rate is more than standard labor rate.Hence variance is unfavorable)
Labor efficiency variance=Standard labor rate*(Standard labor hours required-Actual labor hours worked)
Standard labor rate=$22 per hour
Standard labor hours required=Actual units produced*Labor hours required per unit=(2300*0.10)+(1600*0.45)=950 hours
Actual hours worked=1220 hours
Labor efficiency variance=22*(950-1220)=5940 U
(Actual hours worked is more than the standard hours required.Hence, variance is unfavorable)
Finishing:
Labor rate variance=Actual hours worked*(Statndard labor rate-Actual labor rate)
Actual hours worked=2930 hours
Standard labor rate=$21 per hour
Actual labor rate=Actual direct labor cost/Actual hours=67390/2930=$23 per hour
Labor rate variance=2930*(21-23)=5860 U
(Actual labor rate is more than standard labor rate.Hence variance is unfavorable)
Labor efficiency variance=Standard labor rate*(Standard labor hours required-Actual labor hours worked)
Standard labor rate=$21 per hour
Standard labor hours required=Actual units produced*Labor hours required per unit=(2300*0.80)+(1600*0.90)=3280 hours
Actual hours worked=2930 hours
Labor efficiency variance=21*(3280-2930)=7350 F
(Actual hours worked is less than the standard hours required.Hence, variance is favorable)